### v1

# DeepMIDE: A Multivariate Spatio-Temporal Method for Ultra-Scale Offshore Wind Energy Forecasting

Abstract
To unlock access to stronger winds, the offshore wind industry is advancing with significantly larger and taller wind turbines. This massive upscaling motivates a departure from univariate wind forecasting methods that traditionally focused on a single representative height. To fill this gap, we propose DeepMIDE—a statistical deep learning method which jointly models the offshore wind speeds across space, time, and height. DeepMIDE is formulated as a multi-output integro-difference equation model with a multivariate, nonstationary, and state-dependent kernel characterized by a set of advection vectors that encode the physics of wind field formation and propagation. Embedded within DeepMIDE, an advanced deep learning architecture learns these advection vectors from high-dimensional streams of exogenous weather information, which, along with other parameters, are plugged back into the statistical model for probabilistic multi-height space-time forecasting. Tested on real-world data from future offshore wind energy sites in the Northeastern United States, the wind speed and power forecasts from DeepMIDE are shown to outperform those from prevalent time series, spatio-temporal, and deep learning methods.
Keywords: Forecasting, Integro-difference Equation, Statistical Deep Learning, Spatio-Temporal Modeling, Wind Energy
1 Introduction
Offshore wind energy is poised to play a major role in the worldwide transition towards cleaner and more sustainable energy systems (Musial et al., 2023). To harness stronger winds, the offshore wind industry is advancing with the next-generation offshore wind turbines, which are going to be significantly taller and larger than conventional wind turbines. With hub heights exceeding $140$ meters (m) and rotor diameters as large as $220$m, these ultra-scale turbines are set to become the largest rotating machines on earth, approaching the altitude of many of the world’s tallest landmarks (Gaertner et al., 2020; General Electric, 2021; Zahle et al., 2024). By producing significantly higher per-capita energy output using the same wind resources, considerable reductions in the cost of harnessing wind energy resources can be realized relative to conventional wind turbine designs that typically have had hub heights and rotor diameters of $100$m or less (U.S. Department of Energy, 2021; Shields et al., 2021).
While the massive upscaling of offshore wind turbines unlocks significant benefits related to the “economy of scale,” it also raises considerable challenges pertaining to the optimal operation of those massive-scale assets in a highly dynamic and uncertain environment. Today, the reliable management of offshore wind farms heavily relies on short-term wind speed and power forecasts to mitigate the uncertainty of the intermittent wind resources (Giebel et al., 2016). Those forecasts have been traditionally issued using statistical models that describe the temporal evolution of wind conditions at a single representative height, potentially combining information from multiple geographical locations (Sweeney et al., 2020). By doing so, those models compress the three-dimensional reality of the wind field (space $\times$ time $\times$ height) into a simpler, two-dimensional “world view” (space $\times$ time) which completely overlooks the vertical dimension of the wind field. In this paper, “space” refers specifically to the horizontal dimensions (latitude and longitude), while “height” refers to the vertical dimension (altitude).
While this simplification may be tenable for conventional wind turbines, it is indeed not plausible for the next-generation of offshore wind turbines—there is growing evidence in the wind power engineering literature that larger-scale wind turbines experience a significant influence from vertical wind shear on the turbine’s power production. Hence, the power output of those ultra-scale generators more accurately correlates with the full vertical wind profile swept by the now-much-larger rotor diameter, rather than with a single horizontal slice representing the average hub-height wind speeds (Antoniou et al., 2009; Wagner et al., 2011; Barthelmie et al., 2020). This observation motivates us to seek more sophisticated models for offshore wind forecasting in which the evolution of wind conditions at multiple locations and heights are described in tandem. Our survey of the literature, presented in Section 2, reveals that there is very limited work on multi-height space-time wind forecasting. We believe that, until recently, this has been largely due to the lack of a practical need for such models. Evidently, this is no longer the case with the emergence of ultra-scale wind turbines.
As we argue in this paper, making the leap from single- to multi-height space-time wind forecasting is not a trivial task, primarily due to the complex and non-separable dependencies of the full-fledged wind field across the space-time-height trio. To fill this gap, we propose a multi-output spatio-temporal model with a multivariate, nonstationary, and state-dependent kernel which—unlike existing space-time wind forecasting approaches—is capable of adequately modeling the complex dependencies in local wind speeds across multiple heights. Embedded within the statistical model, an advanced deep learning architecture fully exploits the wealth of exogenous information available to the forecaster (typically in the form of high-dimensional image streams) in order to learn and update key physically meaningful kernel parameters that encode physically relevant information about wind field dynamics. In that way, our proposed approach retains the desirable properties of a statistical model, while embedding a deep learning architecture to act as an internalized “physics extractor” that exploits the large amounts of high-dimensional exogenous information that is difficult to process using standard statistical machinery. Experiments using real-world data from the offshore wind energy areas in the Northeastern United States show that the wind speed and power forecasts from the proposed model are of considerably higher quality than those from prevalent time series, deep learning, and space-time methods.
The remainder of this paper is organized as follows. Section 2 reviews the literature on space-time wind forecasting in the context of wind energy applications. Section 3 presents the real-world data used in this study and its relevance to the rising offshore wind sector in the Northeastern United States. In Section 4, we present the building blocks of the proposed model, followed by Section 5 where forecast evaluations, results, and findings are discussed. Section 6 concludes the paper and highlights future research directions.
2 Literature review
The engineering literature on wind turbine modeling acknowledges the influence of vertical wind speed variations on the performance of large-scale wind turbines (Antoniou et al., 2009; Wagner et al., 2011; Lee et al., 2015; Barthelmie et al., 2020; International Electrotechnical Commission, 2022). This literature primarily focuses on modeling a turbine’s power output, conditional on a fully known wind speed profile.
In practice, however, wind speed—the key determinant of wind power output—is the major source of uncertainty in wind energy harnessing, and is not known to wind farm operators in advance. Thus, to optimally operate an (offshore) wind farm, accurate wind speed and power forecasts are needed at various forecast horizons, ranging from a few minutes up to several days ahead. Those forecasts directly inform critical decisions made by energy producers and grid operators alike, including power production estimation (Zhu et al., 2021; Nasery and Ezzat, 2023), electricity market participation and trading (Pinson, 2013), operations and maintenance scheduling (Papadopoulos et al., 2021, 2024b), economic dispatch and unit commitment (Xie et al., 2013; Barry et al., 2022), among others. Despite the general consensus in the wind engineering community on the importance of considering the vertical wind profile, there are very few studies in the forecasting literature on multi-height wind forecasting (Lin and Liu, 2020; Saxena et al., 2021). Even these studies are only focused on a single location, thereby ignoring the spatial variations and correlations in the wind field.
On the other hand, there is an extensive body of literature on space-time wind forecasting, which focuses on modeling the spatial and temporal variations and correlations, but ignores the vertical dimension of wind fields. Perhaps the most prevalent class of models are geostatistical approaches, with spatio-temporal Gaussian processes (GPs) at their heart. GPs make space-time forecasts that are enabled by a covariance function describing the dependence structure over space and time (Gneiting, 2002; Stein, 2005; Ezzat, 2020). A central challenge in GP-based models is to propose a mathematically permissible covariance function that adequately models the complex dependencies in wind fields. One particulary important aspect of that complexity is to model the effect of time-varying wind advection, which gives rise to an asymmetry in the spatial-temporal correlations. This complexity requires a covariance model that relaxes the convenient, (but unrealistic) assumptions of stationarity, separability, and full space-time symmetry (Gneiting et al., 2006a; Ezzat et al., 2018, 2019).
An alternative paradigm to geostatistical approaches are the so-called dynamic spatio-temporal models (DSTMs) which describe the conditional evolution of a spatial process over time (Wikle and Hooten, 2010). The most prevalent example is the integro-difference equation (IDE) framework, which models the dependence between a spatial process at a future time conditioned on the same process at the present time through an integral operator with a spatial redistribution kernel (Wikle and Cressie, 1999). IDE models have been proposed to predict spatio-temporal cloud movement (Wikle, 2002), precipitation (Ke Xu and Fox, 2005; Liu et al., 2022), aerosols (Calder et al., 2011), and more recently, wind power (Ye and Ezzat, 2024). There are pros and cons for the geostatistical and dynamic space-time approaches, with the former typically regarded as superior surrogate models, whereas the latter are well-suited for extrapolation of highly dynamic processes due to their explicit conditional evolution mechanisms.
Different in principle than the first two statistical approaches (i.e., geostatistical models and DSTMs), deep learning methods have demonstrated significant promise in wind forecasting applications by developing architectures that are tailored to extract complex patterns from spatio-temporal wind data (Wang et al., 2021), including models based on convolutional neural networks (CNNs) (Zhu et al., 2019), recurrent neural networks (RNNs) (Ghaderi et al., 2017), graph neural networks (GNNs) (Khodayar and Wang, 2019; Li et al., 2023), and more recently, transformer models (Bentsen et al., 2023). Despite their demonstrated predictive performance, deep learning still falls short in terms of uncertainty quantification and interpretability. Thus, recent lines of research seek hybrid methods, referred to collectively as statistical deep learning approaches, in which elements or characteristics of a deep learning architecture are embedded within a statistical model to perform certain learning tasks. For example, deep Kalman filters (Krishnan et al., 2015) and deep state-space models (Rangapuram et al., 2018) learn the parameters of a state-space model using multi-layer perceptrons (MLPs), and recurrent neural networks (RNNs), respectively. Deep autoregressive models (DeepAR) entail a sequence of RNNs that are structured in an autoregressive-inspired manner (Salinas et al., 2020), wheras deep IDE models invoke a CNN to learn the IDE redistribution kernel parameters using past process observations (De Bézenac et al., 2019; Zammit-Mangion and Wikle, 2020).
Departing from the existing literature, we propose a multi-output IDE model characterized by a multivariate, nonstationary, and state-dependent redistribution kernel that is tailored to capture the complex dependencies and interactions of local wind fields across the space-time-height trio. To exploit the wealth of exogenous information available to the forecaster (typically in the form of high-dimensional images), we embed an advanced deep learning architecture within the multivariate IDE model in order to learn key physically meaningful kernel parameters that encode physically relevant information about local wind field dynamics. The proposed model is therefore statistical in nature, enabling full probabilistic inference and forecasting across space, time, and height, with the internalized deep learning architecture acting as a “physics extractor,” to learn a low-dimensional set of physically relevant parameters using high-dimensional data that are difficult to process using standard statistical machinery.
Our proposed model departs from the vast majority of space-time wind forecasting models (be it geostatistical, DSTMs, or deep-learning-based) that adopt a two-dimensional world view of the wind field (space $\times$ time) by focusing on a single representative height. It belongs to the class of statistical deep learning models reviewed above, and especially builds on the deep IDE models therein (De Bézenac et al., 2019; Zammit-Mangion and Wikle, 2020), but differs in two key aspects. First, it generalizes the univariate IDE approaches to the multivariate setting to model the complex and time-varying dependencies across space, time, and height. In that way, the proposed model is regarded as the multivariate extension of the univariate IDE model, akin to how multi-output GPs generalize their univariate counterparts. Second, unlike existing deep IDE models which use information that is endogenous to the process (e.g., past observations) to train simple deep learning architectures (typically a CNN), our model exploits the wealth of exogenous information available to the forecaster, typically in the form of high-dimensional image streams. For example, when forecasting wind speeds, one has access to large amounts of high-dimensional data from regional numerical weather models about the meso-scale wind velocities, pressure and temperature systems, etc. that encode rich physical information about the local wind dynamics. To fully unlock access to this exogenous information, we propose to embed a transformer model—an advanced deep learning architecture with evidenced success in computer vision and high-dimensional data mining (Vaswani et al., 2017)—within the multivariate IDE framework as a “physics extractor,” i.e., to learn and update a low-dimensional set of time- and height-varying parameters encoding the physics of local wind field formation and propagation.
3 Data description and analysis
This work is motivated by the ongoing large-scale offshore wind developments in the Northeastern United States, and in particular, the NY/NJ Bight (Bureau of Ocean Energy Management, 2021). Several of those offshore wind projects have already announced plans to install ultra-scale offshore wind turbines. We make use of two sources of data with varying spatial and temporal resolutions: (i) A set of wind speed measurements collected, at multiple heights, by three floating lidar buoys; and (ii) a set of numerical weather predictions (NWPs), of multiple meteorological variables, obtained from a regional numerical weather model. Details of both sets of data are described next.
Figure 1: (a) Locations of the three sites, E05N, E06, and ASOW6, on top of the offshore wind energy areas in the NY/NJ Bight. The background map is generated using the Northeast Ocean Data Portal (Northeast Ocean Data Portal, 2023); (b) Wind rose plot for the wind speeds, averaged over space and height; (c) Time series of wind speeds at all locations (different colors) and heights (different markers), showing strong spatial, temporal, and vertical dependencies, and motivating the need for a joint modeling approach across space, time, and height, as the one advocated in this paper.
3.1 Local observations from the NY/NJ Bight
Wind speeds are measured using three floating lidar buoys, namely E05 Hudson North (E05N), E06 Hudson South (E06) (New York State Energy Development Authority, 2019), and ASOW6 (Atlantic Shores Offshore Wind, 2021). Figure 1(a) shows the locations of the three sites on top of the designated offshore wind energy areas in the NY/NJ Bight. E05N is the farthest location from shore, and is $\sim$77777777 kilometers (km) away from E06. ASOW6 is the southernmost and closest location to shore. E06 and ASOW6 are $\sim$54545454 km apart. The observations are recorded in $10$-min resolution, for a total duration of approximately $8.5$ months spanning the years of 2020 and 2021, at three heights: $100$m, $140$m, and $180$m. The data from ASOW6 had some missing data streaks, which were imputed using another buoy in close proximity. Figure 1(b) shows the wind rose plot of the wind speed data (pooled across all heights and locations), which shows dominant southwesterly winds.
Figure 1(c) shows the multivariate time series of the wind speeds at all locations and heights during a select period in March 2020, clearly demonstrating strong spatial, temporal, and vertical dependencies. Pairwise Pearson correlation coefficients across the three spatial sites (averaged over height) ranged between $0.853$ (between E05 and ASOW6), to $0.928$ (between E06 and ASOW6), suggesting (perhaps unsurprisingly) strong and distant-dependent spatial correlations. Correlations of wind speeds across multiple heights (averaged over space) are estimated at $0.974$ (between $100$m and $180$m), $0.988$ (between $100$m and $140$m), and $0.992$ (between $140$m and $180$m), suggesting strong height dependencies as functions of the vertical separation. Interestingly, height correlations (at the same vertical separation) appear to be stronger at larger altitudes. Height correlations are more pronounced than those over space, further attesting to the need for modeling the multi-height time series jointly, in contrast to the vast majority of existing space-time wind forecasting methods.
3.2 Exogenous information from a regional NWP model
A daily real-time version of the Weather Research and Forecasting (WRF) model, called RU-WRF, generates hourly outputs, at $3$-km resolution, for several meteorological variables, including multi-height wind speeds, sea-surface temperature, humidity, and pressure (Dicopoulos et al., 2021). The model is tailored to the U.S. Mid- and North- Atlantic offshore wind energy areas. It has been independently validated by the National Renewable Energy Laboratory (NREL) (Optis et al., 2020) and has been continuously evaluated and improved since then (Dicopoulos et al., 2021). Its spatial domain spans the coastal region from southern Massachusetts to North Carolina. Figure 2 shows RU-WRF outputs, in the form of spatial weather maps, on a select day and time in June, 2021 on top of the offshore wind energy areas in the NY/NJ Bight.
Figure 2: Spatial weather maps from RU-WRF (wind speed at 100/140/180m, sea-surface temperature, humidity, and pressure) on June 11th, 2021 at 8:00 GMT on top of the offshore wind energy areas (denoted by dashed polygons) in the NY/NJ Bight. Red stars denote the spatial locations where local measurements are available.
3.3 Asymmetry analysis
Due to advection effects, wind fields are known to exhibit strong signs of asymmetry in spatio-temporal dependencies. This means that along-wind dependence (i.e., the dependence between the wind conditions at an upstream location at time $t$ and that at a downstream location at time $t+u$) is expected to be stronger than opposite-wind dependence (i.e., the dependence between the wind conditions at a downstream location at time $t$ and that at an upstream location at time $t+u$) (Gneiting, 2002; Stein, 2005; Ezzat et al., 2018). Thus, adequate models should be able to capture those complex interaction effects. To demonstrate the existence of asymmetry in our data, we first de-trend the time series data at each height and location by fitting a time- and height-specific trend comprising a series of diurnal harmonics, as expressed in (1).
$\small D_{t}^{g}(\mathbf{s})=d_{0}^{g}(\mathbf{s})+d_{1}^{g}(\mathbf{s})\sin\left(\frac{2\pi t}{24}\right)+d_{2}^{g}(\mathbf{s})\cos\left(\frac{2\pi t}{24}\right)+d_{3}^{g}(\mathbf{s})\sin\left(\frac{4\pi t}{24}\right)+d_{4}^{g}(\mathbf{s})\cos\left(\frac{4\pi t}{24}\right),$ (1)
where $D_{t}^{g}(\mathbf{s})$ is the diurnal trend at time $t\in\mathbb{Z}^{+}$, location $\mathbf{s}\in\mathbb{R}^{2}$, and height $g\in\mathbb{Z}^{+}$, whereas $d_{0}^{g}(\mathbf{s}),\ldots,d_{4}^{g}(\mathbf{s})$ are the correspondent regression coefficients. The de-trended data is then used to compute an estimate of asymmetry at the $g$th height, denoted as $a^{g}(\mathbf{s}_{i},\mathbf{s}_{j},u)$, and defined as the difference in spatio-temporal semi-variograms (as a proxy for differences in spatio-temporal covariances). This is expressed in (2).
$a^{g}(\mathbf{s}_{i},\mathbf{s}_{j},u):=\delta^{g}(\mathbf{s}_{i},\mathbf{s}_{j},u)-\delta^{g}(\mathbf{s}_{j},\mathbf{s}_{i},u),$ (2)
where $\mathbf{s}_{i}$ and $\mathbf{s}_{j}$ denote the coordinates of the $i$th and $j$th locations, $u$ is the time lag, and $\delta^{g}(\cdot,\cdot,\cdot)$ is the spatio-temporal semi-variogram at height $g$, which is defined as in (3).
$\delta^{g}\left(\mathbf{s}_{i},\mathbf{s}_{j},u\right)=\frac{1}{2(N-u-1)}\sum_{t=1}^{N-u-1}\left\{r^{g}_{t+u}\left(\mathbf{s}_{i}\right)-r^{g}_{t}\left(\mathbf{s}_{j}\right)\right\}^{2},$ (3)
where $N$ is the total number of data points, and $r^{g}_{t}\left(\mathbf{s}_{j}\right)$ denote the value of the de-trended time series at location $\mathbf{s}_{j}$, time $t$, and height $g$.
Figure 3 shows the average values of $a^{g}(\cdot,\cdot,\cdot)$ versus the time lag $u$ (x-axis), for different heights (different columns), partitioned into two wind speed regimes (different rows), representing weak and strong winds respectively (as defined by a threshold of $8$ m/s). A few key observations can be deduced. First, it is clear that there is noticeable levels of spatial-temporal asymmetry in our data, as evident by the departure from the full-symmetry baseline represented by the horizontal dashed line at 0. This asymmetry is the result of the wind propagation across the prevailing southwesterly wind (recall Figure 1(b)). The asymmetry appears to be maximal at the $\sim$1111-$3$-hour time lag, which aligns with the expected time for wind conditions to propagate across the spatial locations considered. Interestingly, asymmetry appears to be stronger with larger altitudes and stronger winds. This means that changes in wind speed over time and height would be accompanied by changes in the strength of asymmetry. An adequate model would not only need to account for the existence of asymmetry, but also for the time- and height-varying dependence structure driven by wind advection dynamics. Existing asymmetric space-time models typically assume a constant level of asymmetry across height and time (Stein, 2005; Gneiting et al., 2006a), or at best, may model some time-varying advection effects, but overlook the vertical variations in asymmetry levels (Ye et al., 2024).
Figure 3: Asymmetry in space-time dependencies, versus the time lag in hours, for different heights (different columns). The top and bottom rows represent weak and strong wind regimes, respectively. Different colors denote different pairs of locations.
4 Metholodogy
To model the complex dependencies in the local wind fields across space, time, and height, we propose a multi-output deep IDE model with a multivariate, non-stationary, and state-dependent kernel, for which key physically meaningful parameters are learned using an embedded deep learning model. Our approach is dubbed as the Deep Multivariate Integro-Difference Equation model, or in short, DeepMIDE. We start by a brief background to the IDE model, then present the essence and technical details of DeepMIDE.
4.1 Background to the IDE framework
Stochastic IDE models are a class of hierarchical space-time models comprised of a data model and a process model (Cressie and Wikle, 2011). The data model relates a latent state process to a set of observable measurements. In the linear-Gaussian case, it can be written as in (4).
$\mathbf{z}_{t}=\mathbf{h}_{t}\mathbf{y}_{t}+\boldsymbol{\xi}_{t},$ (4)
where $\mathbf{z}_{t}$ is an $m$-dimensional vector of spatial observations at time $t$, while $\mathbf{y}_{t}$ is the correspondent $n$-dimensional vector of the latent state process at time $t$. The matrix $\mathbf{h}_{t}$ is an $m\times n$ observation mapping matrix, whereas $\boldsymbol{\xi}_{t}$ is an $m$-dimensional vector of measurement errors, which are independent of $\mathbf{y}_{t}$. Assuming locally linear dynamics, the process model describes the conditional first-order evolution of the state process through a spatial integral operator, as expressed in (5).
[…]
where ${y}_{t}(\cdot)$ denotes the spatial process on domain $\mathcal{D}_{s}$ at time $t$. The error process $\nu_{t}(\cdot)$ is independent in time but may be correlated in space, whereas $k(\mathbf{s},\mathbf{x};\boldsymbol{\theta})$ is a spatial redistribution kernel characterized by a set of parameters denoted by $\boldsymbol{\theta}$.
Stochastic IDEs have received significant attention in the space-time modeling literature due to their versatility and ability to integrate mechanistic descriptions of physical and domain knowledge. To date, the overwhelming majority of IDE models have been univariate, i.e. they model a single stochastic process over space and time (Wikle, 2002; Xu et al., 2005; Ke Xu and Fox, 2005; Wikle and Holan, 2011; Calder et al., 2011; Liu et al., 2022; Ye and Ezzat, 2024). However, many real-world processes do not exist in isolation; they are influenced by, and interact with, other concomitant processes. Very little attention has been dedicated to multivariate IDE models, and even less so for wind forecasting applications. In Section 4.2, we propose a multivariate IDE to model and forecast the wind speed over space, time, and height, as motivated by large-scale offshore wind turbines.
4.2 A multivariate IDE model for multi-height space-time offshore wind forecasting
Let $Z^{(g)}_{t}(\mathbf{s})$ be the wind speed observation at location $\mathbf{s}$, time $t$, and height $g$. We denote by $\mathbf{Z}^{(g)}_{t}\equiv(Z^{(g)}_{t}(\mathbf{s}_{1}),\ldots,Z^{(g)}_{t}(\mathbf{s}_{m}))^{\top}$ the $m$-dimensional spatial vector of observations at a particular time and height, whereas $\mathbf{Z}_{t}\equiv(\mathbf{Z}^{(1)}_{t},...,\mathbf{Z}^{(G)}_{t})^{\top}$ is the $mG\times 1$ matrix comprising the observations at time $t$, across all locations and heights, such that $G$ is the number of heights where measurements are available. We express the data model for the multivariate IDE as in (6).
$\mathbf{Z}_{t}=\mathbf{H}_{t}\mathbf{Y}_{t}+\boldsymbol{\epsilon}_{t},$ (6)
where $\mathbf{Y}_{t}\equiv(\mathbf{Y}^{(1)}_{t},\ldots,\mathbf{Y}^{(P)}_{t})^{\top}$ is the $nP\times 1$ matrix representing the underlying state process at time $t$, across all locations and heights. The matrix $\mathbf{H}_{t}$ is a $mG\times nP$ observation mapping matrix, whereas $\boldsymbol{\epsilon}_{t}$ denotes the zero-mean measurement error vector, which we assume to be a multivariate Gaussian that is independent in time but not in space, with a spatial covariance matrix $\Sigma_{\boldsymbol{\epsilon}}$, for which the entries are determined using a squared exponential covariance function $c_{\epsilon}(d)=\sigma_{\epsilon}^{2}\exp{(-\frac{1}{2}(\frac{\|d\|^{2}}{\ell^{2}_{\epsilon}})})$, where $d$ is the pairwise spatial distance, while $\sigma_{\epsilon}$ and $\ell_{\epsilon}$ are the variance and length-scale parameters, respectively. The Gaussian assumption for the measurement error is non-problematic after an appropriate Box-Cox transformation which makes the wind speeds nearly Gaussian. For our data, the Box-Cox power parameter is estimated at $0.504$, which is almost equivalent to the square root transformations common for modeling wind speeds in the literature (Haslett and Raftery, 1989; Stein, 2005; Gneiting et al., 2006b).
Considering multivariate spatio-temporal processes representing the offshore wind speeds in a space-height continuum, and discrete time, $\{Y_{t}^{(p)}(\cdot)\}$, then the multivariate version of the stochastic IDE process model can be written as in (7).
[…]
such that $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})$ is the redistribution kernel of the multivariate IDE, which determines the contribution of height $q$ at location $\mathbf{x}$ and time $t-1$ in describing the height $p$ at location $\mathbf{s}$ at the current time $t$. The redistribution kernel is assumed to depend on a set of time-varying parameters, denoted by $\boldsymbol{\theta}_{t}^{(pq)}$. The state error vector at time $t$, across all heights and locations, is denoted by $\boldsymbol{\eta}_{t}=(\boldsymbol{\eta}_{t}(\mathbf{s}_{1}),...,\boldsymbol{\eta}_{t}(\mathbf{s}_{n}))^{\top}$ such that $\boldsymbol{\eta}_{t}(\mathbf{s})\equiv(\eta_{t}^{(1)}(\mathbf{s}),...,\eta_{t}^{(P)}(\mathbf{s}))^{\top}$ is the corresponding error vector at location $\mathbf{s}$ and time $t$, across all $P$ heights. The process error term is also assumed to be a zero-mean Gaussian which is independent in time but not in space, with a spatial covariance matrix $\Sigma_{\boldsymbol{\eta}}$, for which the entries are determined using a squared exponential covariance function $c_{\eta}(d)=\sigma_{\eta}^{2}\exp{(-\frac{1}{2}(\frac{\|d\|^{2}}{\ell^{2}_{\eta}})})$, where $\sigma_{\eta}$ and $\ell_{\eta}$ are the variance and length-scale parameters, respectively.
Modeling the redistribution kernel $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)})$ is perhaps the most pivotal aspect of IDE models. The asymmetry analysis in Section 3.3 demonstrated the need for capturing time- and height-varying advection dynamics. Here, we follow a similar pursuit to the recent work of Salvaña et al. (2023) who introduced asymmetric cross-covariance functions with multiple advections. We leverage their key theorem to propose a multivariate redistribution kernel, as expressed in (8), that is tailored to model the dependence structure in offshore wind fields, considering time- and height-varying advections.
$k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)}\right)=\exp\left(-\frac{\|\mathbf{s}-\mathbf{x}-\boldsymbol{\theta}_{t}^{p}t+\boldsymbol{\theta}_{t}^{q}(t+1)\|^{2}}{\ell_{\mathbbm{1}_{p=q}}^{2}}\right),$ (8)
where $\ell_{\mathbbm{1}_{(p=q)}}$ $\forall$ $p,q\in\{1,\ldots,P\}$ denote a set of height-dependent diffusion parameters, $\mathbbm{1}_{(\cdot)}$ is the indicator function, and the spatial location $\mathbf{s}$, $\mathbf{x}$ in this paper is vectors containing latitude and longitude. The kernel $k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\theta}_{t}^{(pq)}\right)$ is characterized by a set of time- and height-varying advection vectors $\boldsymbol{\theta}_{t}^{(pq)}=\{\boldsymbol{\theta}_{t}^{p},\boldsymbol{\theta}_{t}^{q}\}$, which collectively encode the physics of local wind field formation and propagation, and are assumed to be random in $\mathbb{R}^{2}$.
Collecting all the advection vectors across all heights yields the set $\boldsymbol{\theta}_{t}=\{\boldsymbol{\theta}_{t}^{p}\}_{p=1}^{P}$. Doing so across all time indices yields the parameter set $\boldsymbol{\Theta}=\{\boldsymbol{\theta}_{t}\}_{t=1}^{T}$. This approach, which treats the advection parameters as time- and height-varying, is in line with our local wind dynamics. However, it requires an estimation of a very large number of parameters at each time $t$ and height $p$. Instead, we propose to embed a mapping function, denoted by $\psi(\cdot)$, to represent the parameter set $\boldsymbol{\Theta}$, given streams of exogenous weather maps $\mathcal{X}=\{\mathbf{X}_{t}\}_{t=1}^{T}$ from a regional numerical model, where $\mathbf{X}_{t}\in\mathbb{R}^{C\times\mathcal{W}}$ denotes a set of weather maps at time $t$ representing $C$ exogenous variables, with each map acquired on a bounded rectangle of $\mathbb{R}^{2}$ denoted by $\mathcal{W}$. An example of $\mathbf{X}_{t}$ in the context of our problem are the weather maps shown in Figure 2. Thus, we have $\boldsymbol{\Theta}=\psi(\mathcal{X},\boldsymbol{\Phi})$, where $\boldsymbol{\Phi}$ denote the parameters of the functional mapping, $\psi$. The redistribution kernel in (8) can now be rewritten as in (9), yielding a multivariate, nonstationary, state-dependent redistribution kernel.
$k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)=\exp\left(-\frac{\|\mathbf{s}-\mathbf{x}-\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{p}t+\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{q}(t+1)\|^{2}}{\ell_{\mathbbm{1}_{p=q}}^{2}}\right),$ (9)
where $\psi(\mathbf{X}_{t},\boldsymbol{\Phi})^{p}$ denotes the estimate of the advection vector at time $t$ and height $p$.
By incorporating the state-dependent kernel $k^{(pq)}(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi})$ into the formulation of (7) and approximating the double integral at $n$ spatial locations and $P$ heights as sums, the final form of our latent process model can be expressed as in (10).
$Y_{t}^{(p)}(\mathbf{s})=\sum_{\mathbf{x}=\mathbf{s}_{1}}^{\mathbf{s}_{n}}\sum_{q=1}^{P}k^{(pq)}\left(\mathbf{s},\mathbf{x};t,\boldsymbol{\Phi}\right)Y_{t-1}^{(q)}(\mathbf{x})+\eta_{t}^{(p)}(\mathbf{s}).$ (10)
4.3 Deep learning as a physics extractor
The remaining question is the choice of the mapping function $\psi(\cdot)$. Given the high-dimensional nature of $\mathcal{X}$, we propose a deep learning model to approximate the functional mapping $\psi(\cdot)$ and delineate the complex relationships between the exogenous image streams $\mathcal{X}=\{\mathbf{X}_{t}\}_{t=1}^{T}$ and the advection vectors $\boldsymbol{\Theta}$. In that regard, the embedded deep learning model acts as a “physics extractor,” by learning the advection dynamics of local wind fields from the wealth of exogenous information available to the forecaster. The proposed deep learning architecture in this work fuses AlexNet (a deep CNN-based model) (Krizhevsky et al., 2012) with a custom transformer architecture to fully process the spatial and temporal information in $\mathcal{X}$. Transformer models have shown tremendous promise in computer vision applications (Vaswani et al., 2023) and hence, are well-suited to process the stream of high-dimensional weather maps.
Figure 4 provides a high-level illustration DeepMIDE’s workflow, comprising two phases: an offline training stage ($t\in\{1,\ldots,\tau\}$) and an online training and forecasting stage ($t\in\{\tau+1,\ldots,T+h\}$). In the offline stage, the exogenous weather images $\{\mathbf{X}_{t}\}_{t=1}^{\tau}$ are processed through a sequence of deep CNNs to transform the high-dimensional image data into a condensed vector representation $\{\boldsymbol{\lambda}_{t}\}_{t=1}^{\tau}$. Subsequently, this sequence of vectors $\{\boldsymbol{\lambda}_{t}\}_{t=1}^{\tau}$, representing the compressed spatial features, is fed into a transformer model, which models the temporal and vertical dynamics across the spatial features, ultimately outputting the parameters $\boldsymbol{\Theta}$ that govern the advection dynamics for each corresponding time step. The CNN architecture used is shown to the right of Figure 4, whereas the transformer model architecture, along with more details about the deep learning model, are deferred to the supplementary material document appended to this manuscript.
Figure 4: Illustration of the DeepMIDE framework. During both offline and online training, the inputs to the network are the exogenous image streams $\mathbf{X}_{t}$ (e.g., Pressure, Temperature, etc) (offline $t\in 1,\ldots\tau$; online $t\in{\tau+1,\ldots,T+h}$). In the offline phase, the network parameters $\boldsymbol{\Phi}$ are learned jointly with other statistical parameters $\boldsymbol{\Omega}$ by maximizing the likelihood $\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})$. For the online phase, the deep learning model will be frozen, the network parameters $\boldsymbol{\Phi}$ are fixed and the statistical parameters $\boldsymbol{\Omega}$ will be continuously updated in light of new data. The structure of the deep CNN-based AlexNet is shown to the right.
We denote by $\boldsymbol{\Phi}$ the entire set of deep learning parameters (of both the deep CNNs and the transformer), whereas the remaining statistical parameters of the IDE model are denoted by $\boldsymbol{\Omega}=\{\ell_{\mathbbm{1}_{p=q}},\sigma_{\epsilon}$, $\sigma_{\eta}$, $\ell_{\epsilon}$, $\ell_{\eta}\}$. Given $\mathcal{Z}=\{\mathbf{Z}_{t}\}$ $\forall t$, $\boldsymbol{\Phi}$, and $\boldsymbol{\Omega}$, the likelihood of DeepMIDE can be compactly expressed in closed-form as in (11).
$\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})=\prod_{t}\mathcal{P}(\mathbf{Z}_{t}|\boldsymbol{\Omega},\boldsymbol{\Phi}).$ (11)
The exact form of $\mathcal{P}(\mathbf{Z}_{t}|\boldsymbol{\Omega},\boldsymbol{\Phi})$ is discussed in Section 4.4. For the offline stage, we jointly optimize the full parameter set $\boldsymbol{\Phi}$ and $\boldsymbol{\Omega}$ using stochastic gradient descent (Lecun et al., 1998), with backpropagation used to compute the gradient $\nabla_{\boldsymbol{\Phi},\boldsymbol{\Omega}}\mathcal{L}(\mathcal{Z}|\boldsymbol{\Phi},\boldsymbol{\Omega})$, yielding an optimal set of parameters, denoted by $\boldsymbol{\Phi}^{*}$ and $\boldsymbol{\Omega}^{*}$. For the online stage, the deep learning model will be frozen, i.e., the network parameters $\boldsymbol{\Phi}$ are kept fixed at $\boldsymbol{\Phi}^{*}$, whereas the statistical parameters $\boldsymbol{\Omega}$ will be continuously updated in light of new observations. To issue forecasts, the incoming weather maps are fed directly to the pre-trained deep learning model to predict the advection vector parameters $\{\boldsymbol{\theta}_{t}\}_{t=T+1}^{T+h}$, which are then plugged back into the multivariate IDE to output a set of multi-height space-time forecasts for $h$-steps ahead, denoted by $\hat{\mathbf{Z}}_{T+1},\ldots,\hat{\mathbf{Z}}_{T+h}$. An obvious advantage of the online/offline structure is that the deep learning model needs to be trained only once, thereby significantly reducing the computational burden in the online phase, where the model can quickly apply the pre-learnt network parameters to newly revealed weather maps.
4.4 Inference and forecasting using DeepMIDE
As in univariate IDE models, we can leverage the Kalman Filter (KF) machinery to perform inference on the process and obtain the forecasting and filtering distributions. By writing (10) in matrix format, we have:
$\mathbf{Y}_{t}=\mathbf{K}_{t}\mathbf{Y}_{t-1}+\boldsymbol{\eta}_{t},$ (12)
where $\mathbf{K}_{t}$ comprises all the pairwise evaluations of the redistribution kernel. From there, the forecast distribution conditioned on all past measurements is Gaussian, i.e., $\mathbf{Y}_{t}\mid\mathbf{Z}_{1:t-1}\sim\mathcal{N}\left(\mathbf{Y}_{t\mid t-1},\mathbf{P}_{t\mid t-1}\right)$ such that $\mathbf{Y}_{t\mid t-1}$ and $\mathbf{P}_{t\mid t-1}$ are expressed as in (13).
$\mathbf{Y}_{t\mid t-1}$ $\equiv\mathbb{E}\left(\mathbf{Y}_{t}\mid\mathbf{Z}_{1:t-1}\right)=\mathbf{K}_{t}\mathbf{Y}_{t-1\mid t-1},$ (13)
$\mathbf{P}_{t\mid t-1}$ $\equiv\mathbb{E}[(\mathbf{Y}_{t}-\mathbf{Y}_{t\mid t-1})(\mathbf{Y}_{t}-\mathbf{Y}_{t\mid t-1})^{\top}\mid\mathbf{Z}_{1:t-1}]$
$=\Sigma_{\boldsymbol{\eta}}+\mathbf{K}_{t}\mathbf{P}_{t-1\mid t-1}\mathbf{K}_{t}^{\top}.$
The filtering distribution is also Gaussian $\mathbf{Y}_{t}\mid\mathbf{Z}_{1:t}\sim\mathcal{N}\left(\mathbf{Y}_{t\mid t},\mathbf{P}_{t\mid t}\right)$ such that $\mathbf{Y}_{t\mid t}$ and $\mathbf{P}_{t\mid t}$ are expressed as in (14).
$\mathbf{Y}_{t\mid t}$ $\equiv\mathbf{Y}_{t\mid t-1}+\mathbf{A}_{t}\left(\mathbf{Z}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right),$ (14)
$\mathbf{P}_{t\mid t}$ $\equiv\left(\mathbf{I}-\mathbf{A}_{t}\mathbf{H}_{t}\right)\mathbf{P}_{t\mid t-1},$
$\mathbf{A}_{t}$ $\equiv\mathbf{P}_{t\mid t-1}\mathbf{H}_{t}^{\top}\left(\mathbf{H}_{t}^{\top}\mathbf{P}_{t\mid t-1}\mathbf{H}_{t}+\Sigma_{\boldsymbol{\epsilon}}\right)^{-1}.$
From there, the explicit log-likelihood of DeepMIDE, previously presented in a compressed form in (11), can now be written as in (15).
$\log\mathcal{L}(\mathcal{Z}|\boldsymbol{\Omega},\boldsymbol{\Phi})=$ $\sum_{t}\left(-\frac{mG}{2}\log 2\pi-\frac{1}{2}\log|\mathbf{P}_{t\mid t-1}|\right.$ (15)
$\left.-\frac{1}{2}\left(\mathbf{Z}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right)^{\top}\left(\mathbf{P}_{t\mid t-1}\right)^{-1}\left(\mathbf{Z}_{t}-\mathbf{H}_{t}\mathbf{Y}_{t\mid t-1}\right)\right),$
where ||$\cdot$|| ⋅ | is the matrix determinant. The distribution of the $h$-step latent process forecasts $\hat{\mathbf{Y}}_{T+h}\mid\mathbf{Z}_{1:T}\sim\mathcal{N}(\hat{\mathbf{Y}}_{T+h\mid T+h-1},\hat{\mathbf{P}}_{T+h\mid T+h-1})$ can be expressed as:
$\hat{\mathbf{Y}}_{T+h\mid T+h-1}$ $=\mathbf{K}_{T+h}\hat{\mathbf{Y}}_{T+h-1\mid T+h-2},$ (16)
$\hat{\mathbf{P}}_{T+h\mid T+h-1}$ $=\Sigma_{\boldsymbol{\eta}}+\mathbf{K}_{T+h}\hat{\mathbf{P}}_{T+h-1\mid T+h-2}\mathbf{K}_{T+h}^{\top}.$
The final distribution of the $h$-step forecasts $\hat{\mathbf{Z}}_{T+h}\mid\mathbf{Z}_{1:T}\sim\mathcal{N}(\hat{\mathbf{Z}}_{T+h\mid T+h-1},\hat{\mathbf{C}}_{T+h\mid T+h-1})$ can be expressed as in (17).
$\hat{\mathbf{Z}}_{T+h\mid T+h-1}$ $=\mathbf{H}_{T+h}\hat{\mathbf{Y}}_{T+h\mid T+h-1},$ (17)
$\hat{\mathbf{C}}_{T+h\mid T+h-1}$ $=\Sigma_{\boldsymbol{\epsilon}}+\mathbf{H}_{T+h}\hat{\mathbf{P}}_{T+h\mid T+h-1}\mathbf{H}_{T+h}^{\top}.$
5 Real-world experiments and discussions
We focus on 24-hour forecasts at $10$-min resolution. We train and evaluate DeepMIDE at the three locations where data is available: E05N, E06, and ASOW6. The data spans a total duration of approximately $8.5$ months, so we use six months for offline training, and the remainder is used for online training and forecast evaluation, carried out in a rolling fashion. For each roll, we make $24$-hour forecasts, slide by $24$ hours, and repeat the whole process. This corresponds to a number of rolls $\mathcal{R}=70$, yielding a total of $6$ forecasts/hour $\times$ $24$-hour horizon $\times$ $70$ rolls $\times$ $3$ spatial sites $\times$ $3$ heights = $90{,}720$ testing instances (In this paper, the observation mapping matrix $\mathbf{H}_{t}$ is assumed to be an identity matrix, i.e., $m=n=3,G=P=3$). The offline training takes approximetely $7$ hours using an RTX4090 GPU with $46$ GB RAM, while the online training phase takes approximately 3 minutes.
We compare DeepMIDE against four benchmarks representative of the various forecasting approaches in the wind forecasting literature and practice. All benchmark models are initially trained using the full offline data, and then updated in the online stage using newly revealed observations at each forecast roll.
  • PER: The persistence model is widely regarded as a standard forecasting benchmark, for which the performance is often highly competitive at short-term horizons. PER simply assumes that the observed wind speed at the present time will persist into the future. It is implemented separately for each height and location.
  • ARIMA: The Autoregressive Integrated Moving Average (ARIMA) model is a classical time series approach that considers temporal correlations, but overlooks spatial and vertical dependence. It is implemented for each location and height. All model parameters and coefficients are optimized using the pmdarima package in Python.
  • STGP: The spatio-temporal Gaussian process model is a standard geostatistical approach which considers spatio-temporal correlations through a separable covariance function. The model is trained using the fitrgp package in MATLAB.
  • DeepAR: The Deep Autoregressive model (DeepAR) is a deep statistical method, in which a sequence of RNNs are linked in an autoregressive-based manner. We fit the DeepAR using the GluonTS package in Python, and follow the original suggestion of the authors for setting the parameters (Salinas et al., 2020).
We present the forecasting experiments in Section 5.1, followed by some in-depth discussions about the performance of DeepMIDE.
5.1 Forecasting results
Figure 5(a) shows the spatially averaged mean absolute error (MAE) values for all models plotted against the forecast horizon (in hours) for the three heights considered, namely: $100$m, $140$m, and $180$m. Specifically, the MAE for a representative benchmark $\mathcal{B}$, denoted by $\text{MAE}^{(g)}_{h}(\mathcal{B})$, is calculated as in (18).
$\text{MAE}^{(g)}_{h}(\mathcal{B})=\frac{1}{\mathcal{R}\times m}\sum_{i=1}^{\mathcal{R}}\sum_{j=1}^{m}(|{Z^{(g)}_{T+h,i}(\mathbf{s}_{j})}-{\hat{Z}^{(g)}_{T+h,i}(\mathbf{s}_{j})}|),$ (18)
where ${Z^{(g)}_{T+h,i}(\mathbf{s}_{j})}$ and $\hat{Z}^{(g)}_{T+h,i}(\mathbf{s}_{j})$ denote the actual wind speed and forecasts at location $\mathbf{s}_{j}$ and height $g$ for forecasting horizon $h$ and $i$th forecasting roll, respectively. To provide a more condensed view of the results, Figure 5(b) shows the correspondent boxplots of the percentage reduction in MAE across all locations. The percentage reduction is computed using the average of all absolute errors taken across all forecast rolls and locations for the two methods being compared, as expressed in (19).
$\text{IMP}(\mathcal{B}^{*},\mathcal{B})=100\times(1-\frac{\text{MAE}^{(g)}_{h}(\mathcal{B}^{*})}{\text{MAE}^{(g)}_{h}(\mathcal{B})}),$ (19)
where $\mathcal{B}^{*}$ is the proposed DeepMIDE model, while $\mathcal{B}$ is any other benchmark including PER, ARIMA, STGP, DeepAR. A first glance to Figure 5 shows that, except for few instances, DeepMIDE consistently outperforms all benchmarks across all horizons and heights. DeepMIDE is typically followed by DeepAR as a fairly distant second (average percentage reduction in MAE of $3.26$%). PER and ARIMA perform fairly similarly ($3.91$% and $4.10$% percentage reductions in MAE, respectively). Trailing behind is STGP with an average percent reduction in forecast error of $8.16$% in favor of DeepMIDE.
Figure 5: (a) Average MAE values (aggregated in hourly intervals) versus the forecast horizon for all models, at different heights, namely: $100$m, $140$m, and $180$m. (b) Boxplots of the percentage reduction in MAE, across all rolls and forecast horizon, grouped by height.
Surprisingly, we find that STGP demonstrated weaker performance relative to PER and ARIMA. Upon further investigation, we found that STGP performed fairly well in short-term horizons, but suffered the most in longer-term extrapolation (beyond $\sim$6666 hours ahead), primarily due to the tendency of GP-based models to revert to their mean functions as the spatial and temporal correlations gradually die out. Notably, DeepMIDE achieved the best forecast performance for the first hour (average MAE of $0.660$), compared to an MAE of $0.686$ for ARIMA, $0.687$ for STGP, and $0.689$ for PER. Importantly, we observe that the improvement from DeepMIDE compared to the benchmarks increases as the forecast horizon becomes longer.
We also find that, on average, all models (including DeepMIDE) have slightly higher MAE values at higher altitudes. However, this is likely an artifact of the wind speeds being stronger as the altitude increases. When compared to PER and ARIMA, the improvement of DeepMIDE appears to be fairly consistent across all heights (in the range of $\sim$4444-$5$reduction in MAE). For STGP, the improvement is maximal at the $180$-m height, whereas for DeepAR, the opposite is observed. Thus, we do not necessarily see a clear trend in how DeepMIDE performs across different heights compared to the benchmark models.
Figure 6(a) shows the hourly MAE values $\text{MAE}^{(\mathbf{s})}_{h}(\mathcal{B})$ (averaged over height) plotted against the forecast horizon, across the three spatial sites considered: E05N, E06, and ASOW6, which is calculated as in (20).
$\text{MAE}^{(\mathbf{s})}_{h}(\mathcal{B})=\frac{1}{\mathcal{R}\times G}\sum_{i=1}^{\mathcal{R}}\sum_{g=1}^{G}(|{Z^{(g)}_{T+h,i}(\mathbf{s})}-{\hat{Z}^{(g)}_{T+h,i}(\mathbf{s})}|).$ (20)
Figure 6(b) shows the corresponding boxplots of the percentage MAE reduction. The results are largely similar to those presented in Figure 5. However, this time, there is a clear trend in how DeepMIDE performs across different spatial locations, relative to the baseline models. Specifically, the maximal improvement appears at E06 (in the range of $\sim$5555-$10$reduction in MAE). We postulate that this may be due to E06 being at a central location, in between the two other spatial sites, making the gain from including neighborhood information greater. E06 is followed by E05N ($\sim$3333-$8$reduction in MAE) where its position as a downstream location appears to benefit from the advection dynamics modeled in DeepMIDE. In contrast, ASOW6, being both the southernmost and an upstream location, exhibits the smallest improvement from DeepMIDE ($\sim$2222-$7$reduction in MAE).
Figure 6: Average MAE values (aggregated in hourly intervals) versus the forecast horizon for all models, at different locations, namely: E05N, E06, and ASOW6. (b) Boxplots of the percentage reduction in MAE, across all rolls and forecast horizon, grouped by spatial location.
Figure 7 shows the time series of the one-step wind speed forecasts from DeepMIDE on top of the actual measurements at the hub height of $140$m, during a select time interval in June 2021. The time series of the forecasts show faithful agreement with the actual measurements for all three locations. The corresponding $95$% forecast intervals visually show satisfactory coverage and sharpness. This probabilistic nature of IDE-based models make them well-suited to inform subsequent operational decision-making under uncertainty for offshore wind farms (Pinson, 2013; Papadopoulos et al., 2024a).
Figure 7: Time series of $1$-step ahead forecasts from DeepMIDE on top of the actual wind speed observations at the hub height of $140$m for all three locations ($95$% forecast intervals shown).
Wind speed is the major determinant of wind power, and hence, we would like to illustrate the value of our approach for improving wind power prediction accuracy. To do so, we need to convert the wind speed forecasts into the wind power domain. As there are no operational wind farms in the NY/NJ Bight region yet, we leverage an external set of operational data from a wind farm in the United States to construct a statistical wind power curve which will be used for wind-to-power conversion (Ding, 2019). The power curve is multivariate and it takes two inputs: the hub-height wind speed and the vertical wind shear. The latter reflects the vertical change in wind speed. Let the wind shear be denoted by $S$, then an estimate for wind shear is computed as in (21).
$S=\frac{\log{(Z_{2}/Z_{1})}}{\log{(h_{1}/h_{2})}},$ (21)
where $Z_{2}$ and $Z_{1}$ are the wind speeds measured at heights $h_{1}$ and $h_{2}$, respectively (Lee et al., 2015). For our data, wind speeds are measured at heights of $100$m, $140$m, and $180$m, with $140$m considered as the hub-height. Consequently, two wind shear values are calculated: an above-hub wind shear ($S_{a}$) using the $180$m/$140$m pair, and a below-hub wind shear ($S_{b}$) using the $140$m/$100$m pair. Figure 8(a) visualizes the fitted wind power curve (using a multivariate GP model akin to the one proposed by Golparvar et al. (2021)) with hub height wind speed and above-hub wind shear $S_{a}$. The power output is scaled to the $[0,100]$ interval, where $100$ represents the maximum rated capacity. By applying the constructed power curve, we transform the wind speed observations, along with their forecasts from the four competing models into corresponding wind power values.
Figure 8: (a) Fitted wind power curve with hub height wind speed and above-hub wind shear $S_{a}$, (b) the boxplots of improvement for all forecasting rolls, horizons, locations, and heights in wind power.
Figure 8(b) presents the boxplots of the improvement across all forecasting rolls, horizons, locations, and heights in wind power. DeepMIDE significantly outperforms all its competitors, demonstrating an average improvement of $8.89$% in wind power forecasting compared to other benchmarks. Enhanced accuracy in wind speed forecasts directly contributes to greater improvements in wind power predictions. Theoretically, wind speed is related to wind power through a cubic relationship, and hence, large improvements in forecasting wind speed are expected to yield even greater improvements in predicting wind power.
5.2 Discussion
To further elucidate the inner workings of DeepMIDE, we take a closer look at the advection parameter estimates, which are produced using the deep-learning-based functional mapping. Let the norm of the estimated advection vectors at height $g$ and time $t$ denote as $\Lambda_{t}^{g}=\|\boldsymbol{\theta}_{t}^{g}\|$. Figure 9 shows the time series plot of $\Lambda_{t}^{2}$ at height $140$m (corresponding to height index $g=2$ in our study), clearly illustrating a dynamic pattern over time. We zoom in on two consecutive NWP wind speed maps from 21-Jun-2021 02:00:00 to 03:00:00 (orange rectangle, left) and from 04-Jul-2021 05:00:00 to 06:00:00 (green rectangle, right). During the first time interval, the wind speed maps visually show strong advection effects over time, which appears to be captured by the model as reflected by the relatively large advection parameter predictions in the corresponding time series. In contrast, the second time interval shows significantly weaker advection effects (with the two consecutive images being almost static), with correspondingly smaller predicted values in the advection parameter time series. These two examples demonstrate how the deep learning model is able to translate the information from those maps into physically meaningful estimates of advection. Figure 10 shows the boxplots of all $\{\Lambda_{t}^{g}\}_{t=1}^{T}$, partitioned by wind speed (x-axis) and height (different colors), which clearly shows that the values of the predicted advection parameters increase at greater heights and stronger winds—a conclusion which we have observed in our preliminary data analysis of Section 3.3. This further confirms the ability of DeepMIDE to capture real-world aspects of the data.
Figure 9: Time series plot showing the norm of the predicted advection parameters at height $140$m, alongside spatial wind speed images from the numerical model at three consecutive time steps, for two different time intervals: 21-Jun-2021 02:00:00 to 03:00:00 (orange rectangle, left) and from 04-Jul-2021 05:00:00 to 06:00:00 (green rectangle, right).
Figure 10: Boxplot showing the norm of the predicted advection vectors $\{\Lambda_{t}^{g}\}_{t=1}^{T}$. for different heights and wind speeds.
6 Conclusions
As the upscaling of offshore wind turbines to unprecedented heights continues, multi-height wind speed forecasts across the now-much-larger rotor area of those ultra-scale generators are now needed. In this work, we propose the first attempt for multi-height spatio-temporal wind energy forecasting. Our model, dubbed DeepMIDE, is formulated as a multivariate integro-difference equation model, which fully embraces the full vertical wind profile by incorporating spatial, temporal, and height dependencies, in order to reflect the true complexity of wind interactions coincident with ultra-scale offshore wind turbines. An embedded deep learning architecture directly acts on streams of high-dimensional exogenous weather information in order to encode the physics of local wind field formation in a low-dimensional set of information-rich advection vectors. Those vectors, along with other parameters, are then plugged back into the statistical model for probabilistic space-time forecasting. The proposed approach marries the probabilistic rigor and parsimony of the multivariate integro-difference equation framework, with the representative power of the advanced deep transformer architecture which are used as a “physics extractor” within the statistical model. Evaluated using real-world data from the U.S. Mid Atlantic, our model demonstrates significant improvements in both wind speed and power forecasting compared with prevalent forecasting benchmarks.
Future research will investigate more sophisticated modeling structures for the latent process model within the MIDE framework. This exploration may include handling non-Gaussian data and non-linear dependence, which would obviate the need for data transformations or linear assumptions, but would likely complicate the parameter inference process.
Online Supplementary Materials
The supplementary materials appended to this manuscript include: (1) Details of the architecture and implementation of the deep learning model shown in Figure 4; (2) Access to codes and data used to reproduce the results of DeepMIDE.
Disclosure Statement
The authors report there are no competing interests to declare.